# On the cardinality of measures of maximal relative entropy for smooth skew products

### v2

@@ Abstract.
Let $\Omega$ and $M$ be compact smooth manifolds and let $\Theta:\Omega\times M\to\Omega\times M$ be a $\mathcal{C}^{1+\alpha}$ skew-product diffeomorphism over an Axiom A base. We show that $\Theta$ has at most countably many ergodic hyperbolic measures of maximal relative entropy. When $\dim M=2$, if $\Theta$ has positive relative topological entropy, then $\Theta$ has at most countably many ergodic measures of maximal relative entropy.
@@ 1. Introduction and Main results
among all $\pi$-lifts $\eta$ of $\nu$. This relative viewpoint is captured by the relativised variational principle of Ledrappier–Walters and in Walters’ compensation function, extending the notion of relative entropy to relative equilibrium states [30, 52].
The general picture of measures of maximal relative entropy and relative equilibrium states is well advanced in symbolic dynamics, particularly for factor maps between shifts where $X$ is an irreducible subshift of finite type. Petersen–Quas–Shin proved that for each ergodic base measure $\nu$ on $Y$ there are only finitely many ergodic measures of maximal relative entropy on $X$, and they established explicit bounds on this number [44]. These bounds were improved in [2], introducing the notion of class degree. A subsequent work extended these advances to relative equilibrium states in [1]. A structure theorem for infinite-to-one factor codes $\pi\colon X\to Y$ between irreducible subshifts of finite type was proved in [53]: that there exists a sofic shift $Y_{1}$, such that $\pi=\pi_{2}\circ\pi_{1}$. This allowed Yoo to show that if $\nu$ is an equilibrium state on $Y$ for a sufficiently regular potential, then there is a unique lift on $X$ that is a relative equilibrium state [53].
Measures of relative maximal entropy and relative equilibrium states naturally arise in random dynamical systems. There, they are usually called quenched relative equilibrium states. Consider a probability-preserving base $(\Omega,\theta,\mathbb{P})$ and the skew-product map
@@
Foundational results of Kifer established existence and uniqueness of relative equilibrium states for random uniformly expanding transformations and identified quenched relative pressure with the principal Lyapunov exponent of the random transfer-operator cocycle [28]. Bogenschütz developed parallel notions of metric entropy and pressure for random dynamical systems and proved a variational principle, in the framework commonly used nowadays [12]. Thermodynamic formalism for random systems has since been developed across a variety of settings: random symbolic dynamics [13, 22, 20, 49, 48]; smooth expanding maps [28, 38]; non-uniformly smooth expanding maps [5, 50]; and non-uniformly hyperbolic random interval maps (closed and open, with discontinuities) [8, 6, 7]; see also the references therein.
@@
Although the quenched relative thermodynamic formalism theory for random dynamical systems is extensive, results for systems generated by diffeomorphisms remain scarce outside the SRB setting (for SRB results, see [29, 11, 21, 3, 36] and the references therein). In particular, except in settings where Gibbs-type bounds are available, there is no general method for bounding the number of measures of relative maximal entropy, let alone the number of relative equilibrium states, in these settings. In this paper, we study how many measures of relative maximal entropy a skew product $\Theta(\omega,x)=(\theta(\omega),T_{\omega}(x))$ over an Axiom A base may admit. This question is part of an ongoing program to quantify the number of measures of maximal entropy across classes of systems.
@@
In the setting of smooth diffeomorphisms, Newhouse proved that $\mathcal{C}^{\infty}$ diffeomorphisms on compact smooth manifolds always admit a measure of maximal entropy [39]. Buzzi later conjectured that for surface diffeomorphisms with positive topological entropy, there are at most countably many ergodic measures of maximal entropy [16, Conjecture 1]. Sarig’s groundbreaking symbolic coding for $C^{1+\alpha}$ surface diffeomorphisms confirmed this countability and also settled a conjecture of Katok on periodic-orbit growth [45]. This result was recently sharpened by Buzzi–Crovisier–Sarig, proving that $\mathcal{C}^{\infty}$ surface diffeomorphisms with positive entropy have only finitely many ergodic measures of maximal entropy, and exactly one in the transitive case, [18], which answered another of Buzzi’s conjectures [16, Conjecture 2].
@@
In higher dimensions, Ben Ovadia extended Sarig’s coding to non-uniformly hyperbolic diffeomorphisms on arbitrary manifolds, yielding countability results for hyperbolic measures of maximal entropy [10]. More recently, Buzzi–Crovisier–Sarig established the strong positive recurrence (SPR) property, and via SPR obtained strong statistical properties for the associated invariant measures, including exponential decay of correlations and an almost-sure invariance principle [19]. Related results are known for two-dimensional fibre skew products over hyperbolic bases; see Marín–Poletti–Veiga [37]. Analogous results are available for flows [35, 17, 33] and for discontinuous or maps with singularities [32, 31, 4, 34].
@@
Our contribution is a relative analogue of these countable cardinality results. For $\mathcal{C}^{1+\alpha}$ skew–product diffeomorphisms $\Theta(\omega,x)=(\theta\omega,T_{\omega}(x))$ over an Axiom A base $(\Omega,\theta)$ and for any ergodic base-invariant measure $\mathbb{P}$, we prove that there are only countably many ergodic hyperbolic $\Theta$-invariant measures of maximal entropy relative to $\mathbb{P}$. In particular, when the fibre $M$ is a surface and $h_{\mathrm{top}}(\Theta\mid\nu)>0$, the set of ergodic $\mathbb{P}$-relative measures of maximal entropy is countable. The proof refines the symbolic coding methods of [45, 19] to the skew-product setting and combines them with the finiteness techniques for factor maps due to Petersen–Quas–Shin [44]. We expect the coding-and-factor approach developed here to be useful in random dynamics, and to provide a natural quenched perspective and new tools for constructing and counting relative equilibrium states.
@@
The paper is organised as follows. The remainder of Section 1 states the main results and presents an application. Section 2 recalls background from Pesin theory and reviews homoclinic classes. In Section 3 we construct a natural symbolic coding for the skew product $\Theta$ that preserves the base dynamics. Section 4 generalises the results of [44] to countable-state Markov shifts. Finally, Section 5 contains the proofs of the main results, Theorem A and Corollary B build upon the results established in Sections 3 and 4.
@@ 1.1. Measures of maximal relative entropy
This paper concerns the amount of $\mathbb{P}$-relative entropy that $\Theta$ may admit. Below, we define the metric entropy via the Brin–Katok formula [15] since it is the shortest way to define it (see also [51, Chapter 4]). We also define the concept of $\mathbb{P}$-relative entropy for skew products via an extension of the Brin–Katok formula (see [55]).
@@
Let $\Theta$ be as (1) and $\mu$ be a $\Theta$-invariant probability with $(\mathrm{proj}_{\Omega})_{*}\mu=\mathbb{P}$, and disintegrate $\mu(\mathrm{d}\omega,\mathrm{d}x)=\mu_{\omega}(\mathrm{d}x)\,\mathbb{P}(\mathrm{d}\omega).$ For $\omega\in\Omega$, $x\in M$, $\varepsilon>0$, $n\in\mathbb{N}$, write
@@
We recall that if $h_{\mathbb{P}}(\theta)<\infty$, which will always be the case in this paper, we have that $h_{\mu}(\Theta\mid\mathbb{P})=h_{\mu}(\Theta)-h_{\mathbb{P}}(\theta)$ (see [25, Chapter 5, below equation 1.1.5]).
@@
To formulate our results, we impose additional assumptions on the skew-product base $\theta:\Omega\to\Omega$ and the measure $\mathbb{P}$, namely, we assume $\theta$ is an Axiom A diffeomorphism and $\mathbb{P}$ is an ergodic $\theta$-invariant probability measure.
@@
Let $\Omega$ be a compact smooth manifold and $\theta:\Omega\to\Omega$ be $\mathcal{C}^{1}$ diffeomorphism. We say that $\theta$ is an Axiom A transformation if
  • (i)
    $\overline{\{p\in\Omega;\ \text{$p$ is a $\theta$-periodic point}\}}=NW(\theta)$, where
    $NW(\theta):=\left\{\omega\in\Omega\,\middle|\,\begin{aligned} &\text{for each open set }U\subset\Omega\text{ containing }\omega,\\ &\#\{n\in\mathbb{N};\theta^{n}(U)\cap U\neq\emptyset\}=\#\{n\in\mathbb{N};\ \theta^{-n}(U)\cap U\neq\emptyset\}=\infty\end{aligned}\right\};$
    and
  • (ii)
    the non-wandering set $NW(\theta)$ is a uniformly hyperbolic set, i.e. there exists $\chi>0$ and a direct sum $T_{\omega}\Omega=E^{s}(\omega)\oplus E^{u}(\omega)$ for all $\omega\in NW(\theta),$ and a uniform number $K>0$ such that for any $k\geq 0,$ and $\omega^{\prime}\in NW(\theta)$,
    $\mathrm{max}\left(\left\|\left.\mathrm{D}\theta^{k}\right|_{E^{s}(\omega^{\prime})}\right\|,\left\|\left.\mathrm{D}\theta^{-k}\right|_{E^{s}(\omega^{\prime})}\right\|\right)\leq Ke^{-\chi k}.$
@@ 1.2. Main results
Below, we state our main results.
@@
be a $\mathcal{C}^{1+\alpha}$ diffeomorphism, for some $\alpha>0$. Assume that $\theta:\Omega\to\Omega$ is an Axiom $A$ diffeomorphism and $\mathbb{P}$ an ergodic probability measure. Then, $\Theta$ admits at most countably many ergodic hyperbolic measures of maximal $\mathbb{P}$-relative entropy.
Theorem A is proved in Section 5.
We remark that Theorem A is a generalisation of the classical non-skew-product setting. Let $T:M\to M$ be a $\mathcal{C}^{1+\alpha}$ map, $\theta:\Omega\to\Omega$ be any smooth Axiom $A$ map, and consider the product map $\Theta(\omega,x)=(\theta(\omega),T(x))$. Fix a base point $\omega_{p}\in\Omega$ with $\theta(\omega_{p})=\omega_{p}$ and let $\mathbb{P}=\delta_{\omega_{p}}$. Then any $\Theta$-invariant measure $\mu$ satisfying $(\mathrm{proj}_{\Omega})_{*}\mu=\delta_{\omega_{p}}$ has the form $\mu=\delta_{\omega_{p}}\otimes\mu^{\prime}$, where $\mu^{\prime}$ is $T$-invariant on $M$, and
$h_{\mu}\big(\Theta\,\big|\,\delta_{\omega_{p}}\big)=h_{\mu^{\prime}}(T).$
Hence, measures of maximal $\mathbb{P}$-relative entropy for $\Theta$ correspond exactly to measures of maximal entropy for $T$. In particular, Theorem A yields the usual countability statement for ergodic hyperbolic measures of maximal entropy in the non–skew-product setting in conformance with [10, Theorem 0.5].
@@
Moreover, for fibre maps in dimension $n\geq 3$, the countability conclusion for hyperbolic measures of maximal entropy in Theorem A is sharp, i.e. there exists a $\mathcal{C}^{\infty}$ skew-product diffeomorphism $\Theta:\Omega\times\mathbb{T}^{n}\to\Omega\times\mathbb{T}^{n}$ satisfying the hypotheses of Theorem A which admits infinitely many ergodic hyperbolic measures of $\mathbb{P}$-relative maximal entropy. As before, take $\mathbb{P}=\delta_{p}$, where $p\in\Omega$ is a fixed point of the base dynamics and set $\Theta(\omega,x)=(\theta(\omega),T(x))$, where $T:\mathbb{T}^{n}\to\mathbb{T}^{n}$ admits infinitely many hyperbolic measures of maximal entropy. One possible construction of $T$ is as follows.
@@
where $A\in\mathrm{SL}(n-1,\mathbb{Z})$ is an Anosov toral automorphism, and $F\colon\mathbb{S}^{1}\to\mathbb{S}^{1}$ is the function
@@
Since circle diffeomorphisms satisfy $h_{\mathrm{top}}(F)=0$, we obtain that $h_{\mathrm{top}}(T)=h_{\mathrm{top}}(A)$. If $m$ is the (unique) measure of maximal entropy for $A$ (in fact, the Lebesgue measure on $\mathbb{T}^{n-1}$), then for each $j\in\mathbb{N}_{>1}$ the product $m_{j}\coloneqq m\otimes\delta_{1/j}$ is ergodic, hyperbolic, and satisfies $h_{m_{j}}(T)=h_{m}(A)=h_{\mathrm{top}}(T)$. Thus $T$ admits infinitely many ergodic hyperbolic measures of maximal entropy.
@@
In dimension two, it may be possible to strengthen the conclusion of “at most countably many measures of maximal $\mathbb{P}$-relative entropy” of Corollary B to “at most finitely many measures of maximal $\mathbb{P}$-relative entropy”, in analogy with the classical non-skew-product case [18, Corollary 1.4]. Another natural future direction is to derive a criterion for uniqueness. In [18] (see also [19]) it is shown that each homoclinic class supports a unique measure of maximal entropy. In particular, any smooth, transitive surface system with positive entropy has a unique measure of maximal entropy. The analogous statement for measures of relative maximal entropy remains unclear.
It is worth mentioning that, as a consequence of a celebrated result due to Newhouse [39], $\mathcal{C}^{\infty}$ skew products always admit a measure of maximal $\mathbb{P}$-relative entropy.

### v3

@@ Abstract.
Let $\Omega$ and $M$ be compact smooth manifolds and let $\Theta:\Omega\times M\to\Omega\times M$ be a $\mathcal{C}^{1+\alpha}$ skew-product diffeomorphism over a transitive Anosov base. We show that $\Theta$ has at most countably many ergodic hyperbolic measures of maximal relative entropy. When $\dim M=2$, if $\Theta$ has positive relative topological entropy, then $\Theta$ has at most countably many ergodic measures of maximal relative entropy.
@@ 1. Introduction and main results
among all $\pi$-lifts $\eta$ of $\nu$. This relative viewpoint is captured by the relativised variational principle of Ledrappier–Walters and in Walters’ compensation function, extending the notion of relative entropy to relative equilibrium states [LedrappierWalters77, Walters86].
The general picture of measures of maximal relative entropy and relative equilibrium states is well advanced in symbolic dynamics, particularly for factor maps between shifts where $X$ is an irreducible subshift of finite type. Petersen–Quas–Shin proved that for each ergodic base measure $\nu$ on $Y$ there are only finitely many ergodic measures of maximal relative entropy on $X$, and they established explicit bounds on this number [Quas]. These bounds were improved in [AllahbakhshiQuas12], introducing the notion of class degree. A subsequent work extended these advances to relative equilibrium states in [AllahbakhshiAntonioliYoo19]. A structure theorem for infinite-to-one factor codes $\pi\colon X\to Y$ between irreducible subshifts of finite type was proved in [YooDecomp17]: that there exists a sofic shift $Y_{1}$, such that $\pi=\pi_{2}\circ\pi_{1}$. This allowed Yoo to show that if $\nu$ is an equilibrium state on $Y$ for a sufficiently regular potential, then there is a unique lift on $X$ that is a relative equilibrium state [YooDecomp17].
Measures of maximal relative entropy and relative equilibrium states naturally arise in random dynamical systems. There, they are usually called quenched relative equilibrium states. Consider a probability-preserving base $(\Omega,\theta,\mathbb{P})$ and the skew-product map
@@
Foundational results of Kifer established existence and uniqueness of relative equilibrium states for random uniformly expanding transformations and identified quenched relative pressure with the principal Lyapunov exponent of the random transfer-operator cocycle [Kifer92]. Bogenschütz developed parallel notions of metric entropy and pressure for random dynamical systems and proved a variational principle, in the framework commonly used nowadays [Bogenschutz92]. Thermodynamic formalism for random systems has since been developed across a variety of settings: random symbolic dynamics [BogenschutzGundlach1995ETDS, GundlachKifer2000DCDS, DenkerKiferStadlbauer2008DCDS, Stadlbauer2010StochDyn, Stadlbauer2017ETDS]; smooth expanding maps [Kifer92, MSU11]; non-uniformly smooth expanding maps [ArbietoMatheusOliveira03, StadlbauerSuzukiVarandas2021CMP]; and non-uniformly hyperbolic random interval maps (closed and open, with discontinuities) [AtnipFroylandGonzalezTokmanVaienti2021CMP, AtnipFroylandGonzalezTokmanVaienti2023ETDS, AtnipFroylandGonzalezTokmanVaienti2024DM]; see also the references therein.
@@
Although the quenched relative thermodynamic formalism theory for random dynamical systems is extensive, results for systems generated by diffeomorphisms remain scarce outside the SRB setting (for SRB results, see [Ledrappier-Young, BlumenthalYoung2019, DragicevicFroylandGonzalezTokmanVaienti2020TAMS, Alves2023, Liu2024] and the references therein). In particular, except in settings where Gibbs-type bounds are available, there is no general method for bounding the number of measures of maximal relative entropy, let alone the number of relative equilibrium states, in these settings. In this paper, we study how many measures of maximal relative entropy a skew product $\Theta(\omega,x)=(\theta(\omega),T_{\omega}(x))$ over an Anosov base may admit. This question is part of an ongoing program to quantify the number of measures of maximal entropy across classes of systems.
@@
In the setting of smooth diffeomorphisms, Newhouse proved that $\mathcal{C}^{\infty}$ diffeomorphisms on compact smooth manifolds always admit a measure of maximal entropy [Newhouse]. Buzzi later conjectured that for surface diffeomorphisms with positive topological entropy, there are at most countably many ergodic measures of maximal entropy [Buzzi1, Conjecture 1]. Sarig’s groundbreaking symbolic coding for $\mathcal{C}^{1+\alpha}$ surface diffeomorphisms confirmed this countability and also settled a conjecture of Katok on periodic-orbit growth [Sarig2013]. This result was recently sharpened by Buzzi–Crovisier–Sarig, proving that $\mathcal{C}^{\infty}$ surface diffeomorphisms with positive entropy have only finitely many ergodic measures of maximal entropy, and exactly one in the transitive case, [BCS], which answered another of Buzzi’s conjectures [Buzzi1, Conjecture 2].
@@
In higher dimensions, Ben Ovadia extended Sarig’s coding to non-uniformly hyperbolic diffeomorphisms on arbitrary manifolds, yielding countability results for hyperbolic measures of maximal entropy [BenOvadia2018]. More recently, Buzzi–Crovisier–Sarig established the strong positive recurrence (SPR) property, and via SPR obtained strong statistical properties for the associated invariant measures, including exponential decay of correlations and an almost-sure invariance principle [buzzi2025strongpositiverecurrenceexponential]. Related results are known for two-dimensional fibre skew products over hyperbolic bases; see Marín–Poletti–Veiga [marin2025exponentialmixingmeasuresmaximal]. Analogous results are available for flows [Lima, Lima3, lima2025symbolicdynamicsnonuniformlyhyperbolic] and discontinuous maps or maps with singularities [YuriCarlos, Yuri2, Poletti1, lima2024measuresmaximalentropynonuniformly].
@@
Our contribution is a relative analogue of these countable cardinality results. For $\mathcal{C}^{1+\alpha}$ skew–product diffeomorphisms $\Theta(\omega,x)=(\theta(\omega),T_{\omega}(x))$ over a transitive Anosov base $(\Omega,\theta)$ and for any ergodic $\theta$-invariant measure $\mathbb{P}$ with full support, we prove that there are only countably many hyperbolic ergodic $\Theta$-invariant measures of maximal $\mathbb{P}$-relative entropy. In particular, when the fibre $M$ is a surface and $h_{\mathrm{top}}(\Theta\mid\mathbb{P})>0$, the set of ergodic $\mathbb{P}$-relative measures of maximal entropy is countable. The proof refines the symbolic coding methods of [Sarig2013, buzzi2025strongpositiverecurrenceexponential] to the skew-product setting and combines them with the finiteness techniques for factor maps due to Petersen–Quas–Shin [Quas]. We expect the coding-and-factor approach developed here to be useful in random dynamics, and to provide a natural quenched perspective and new tools for constructing and counting relative equilibrium states.
@@
The paper is organised as follows. The remainder of Section 1 states the main results and presents an application. Section 2 recalls background from Pesin theory and reviews homoclinic classes. In Section 3 we construct a natural symbolic coding for the skew product $\Theta$ that preserves the base dynamics. Section 4 generalises the results of [Quas] to countable-state Markov shifts. Finally, Section 5 contains the proofs of the main results, Theorem A and Corollary B build upon the results established in Sections 3 and 4.
@@ 1.1. Measures of maximal relative entropy
This paper concerns the amount of $\mathbb{P}$-relative entropy that $\Theta$ may admit. Below, we define the metric entropy via the Brin–Katok formula [BrinKatok1983LocalEntropy] since it is the shortest way to define it (see also [WaltersBook, Chapter 4]). We also define the concept of $\mathbb{P}$-relative entropy for skew products via an extension of the Brin–Katok formula (see [Yujun]).
@@
Let $\Theta$ be as (1) and $\mu$ be a $\Theta$-invariant probability measure with $(\mathrm{proj}_{\Omega})_{*}\mu=\mathbb{P}$, and disintegrate $\mu(\mathrm{d}\omega,\mathrm{d}x)=\mu_{\omega}(\mathrm{d}x)\,\mathbb{P}(\mathrm{d}\omega).$ For $\omega\in\Omega$, $x\in M$, $\varepsilon>0$, $n\in\mathbb{N}$, write
@@
We recall that if $h_{\mathbb{P}}(\theta)<\infty$, which will always be the case in this paper, we have that $h_{\mu}(\Theta\mid\mathbb{P})=h_{\mu}(\Theta)-h_{\mathbb{P}}(\theta)$ (see [KatokKifer, Chapter 5, below equation 1.1.5]).
@@
To formulate our results, we impose additional assumptions on the skew-product base $\theta:\Omega\to\Omega$ and the measure $\mathbb{P}$, namely, we assume $\theta$ is a transitive Anosov diffeomorphism and $\mathbb{P}$ is a $\theta$-invariant probability measure of full support, i.e.
$\mathrm{supp}\,\mathbb{P}:=\{\omega\in\Omega;\,\mathbb{P}[U]>0\ \text{for every open neighbourhood $U$ of $\omega$}\}=\Omega$
@@
Let $\Omega$ be a compact connected smooth manifold and let $\theta:\Omega\to\Omega$ be a $\mathcal{C}^{1+\alpha}$ diffeomorphism. We say that $\theta$ is Anosov if there exist a continuous $\mathrm{D}\theta$-invariant splitting $T_{\omega}\Omega=E^{s}(\omega)\oplus E^{u}(\omega),$ and constants $K\geq 1$, $\chi>0$ such that for all $k\geq 0$,
$\mathrm{max}\left(\left\|\left.\mathrm{D}\theta^{k}\right|_{E^{s}(\omega^{\prime})}\right\|,\left\|\left.\mathrm{D}\theta^{-k}\right|_{E^{u}(\omega^{\prime})}\right\|\right)\leq Ke^{-\chi k}.$
We say that $\theta$ is a transitive Anosov diffeomorphism if $\theta$ is Anosov and $\theta$ is topologically transitive, i.e. there exists $\omega\in\Omega$ with $\overline{\{\theta^{k}(\omega):k\in\mathbb{Z}\}}=\Omega$.
@@ 1.2. Main results
Below we state our main results. The definition of hyperbolic measure used here is recalled in Definition 2.1.
@@
be a $\mathcal{C}^{1+\alpha}$ diffeomorphism, for some $\alpha>0$. Assume that $\theta:\Omega\to\Omega$ is a transitive Anosov diffeomorphism and $\mathbb{P}$ is a $\theta$-invariant ergodic probability measure with full support. Then, $\Theta$ admits at most countably many hyperbolic ergodic measures of maximal $\mathbb{P}$-relative entropy.
Theorem A is proved in Section 5. We mention that many natural measures satisfy the full-support assumption; for example, since $\theta$ is a transitive Anosov diffeomorphism, every ergodic equilibrium state for a Hölder potential has full support (see [BowenGibbs]).
We remark that Theorem A is a generalisation of the classical non-skew-product setting. Let $T:M\to M$ be a $\mathcal{C}^{1+\alpha}$ map, and $\theta:\mathbb{T}^{2}\to\mathbb{T}^{2}$ be Arnold’s cat map
$\theta\left(\omega_{1},\omega_{2}\right)=\left(2\omega_{1}+\omega_{2},\,\omega_{1}+\omega_{2}\right)\ \mathrm{mod}\ 1.$
Consider the product map $\Theta(\omega,x)=(\theta(\omega),T(x))$ and let $\mathbb{P}$ be Lebesgue measure on $\mathbb{T}^{2}$ (which is $\theta$-ergodic). Then any $\Theta$-invariant measure $\mu$ satisfying $(\mathrm{proj}_{\Omega})_{*}\mu=\mathbb{P}$ has the form $\mu=\mathbb{P}\otimes\mu^{\prime}$, where $\mu^{\prime}$ is $T$-invariant on $M$, and
$h_{\mu}\big(\Theta\,\big|\,\mathbb{P}\big)=h_{\mu^{\prime}}(T).$
Hence, measures of maximal $\mathbb{P}$-relative entropy for $\Theta$ correspond exactly to measures of maximal entropy for $T$. In particular, Theorem A yields the usual countability statement for ergodic hyperbolic measures of maximal entropy in the non–skew-product setting in conformance with [BenOvadia2018, Theorem 0.5].
@@
Moreover, for fibre maps in dimension $n\geq 3$, the countability conclusion for hyperbolic measures of maximal entropy in Theorem A is sharp, i.e. there exists a $\mathcal{C}^{\infty}$ skew-product diffeomorphism $\Theta:\Omega\times\mathbb{T}^{n}\to\Omega\times\mathbb{T}^{n}$ satisfying the hypotheses of Theorem A which admits infinitely many hyperbolic ergodic measures of maximal $\mathbb{P}$-relative entropy. As before, take $\mathbb{P}$ as the Lebesgue measure in $\mathbb{T}^{2}$ and $\theta:\mathbb{T}^{2}\to\mathbb{T}^{2}$ as Arnold’s cat map, and set $\Theta(\omega,x)=(\theta(\omega),T(x))$, where $T:\mathbb{T}^{n}\to\mathbb{T}^{n}$ admits infinitely many hyperbolic measures of maximal entropy. One possible construction of $T$ is as follows.
@@
where $A\in\mathrm{SL}(n-1,\mathbb{Z})$ is a transitive Anosov toral automorphism, and $F\colon\mathbb{S}^{1}\to\mathbb{S}^{1}$ is the function
@@
Since circle diffeomorphisms satisfy $h_{\mathrm{top}}(F)=0$, we obtain that $h_{\mathrm{top}}(T)=h_{\mathrm{top}}(A)$. If $m$ is the (unique) measure of maximal entropy for $A$ (in fact, the Lebesgue measure on $\mathbb{T}^{n-1}$), then for each $j\in\mathbb{N}_{>1}$ the product $m_{j}\coloneqq m\otimes\delta_{1/j}$ is ergodic, hyperbolic, and satisfies $h_{m_{j}}(T)=h_{m}(A)=h_{\mathrm{top}}(T)$. Thus $T$ admits infinitely many hyperbolic ergodic measures of maximal entropy.
@@
In dimension two, it may be possible to strengthen the conclusion of “at most countably many measures of maximal $\mathbb{P}$-relative entropy” of Corollary B to “at most finitely many measures of maximal $\mathbb{P}$-relative entropy”, in analogy with the classical non-skew-product case [BCS, Corollary 1.4]. Another natural future direction is to derive a criterion for uniqueness. In [BCS] (see also [buzzi2025strongpositiverecurrenceexponential]) it is shown that each homoclinic class supports a unique measure of maximal entropy. In particular, any smooth, transitive surface system with positive entropy has a unique measure of maximal entropy. The analogous statement for measures of maximal relative entropy remains unclear. Finally, we remark that the Anosov hypothesis may extend to Axiom A, but the symbolic system for $\Theta$ used in Theorem 3.6 would require a different construction.
It is worth mentioning that, as a consequence of a celebrated result due to Newhouse [Newhouse], $\mathcal{C}^{\infty}$ skew products always admit a measure of maximal $\mathbb{P}$-relative entropy.